OpenStudy (daniarrianna):

something (7kg) is dropped out of a building, falling for 6 seconds. How much work did gravity apply? how much power did gravity apply?

OpenStudy (hoot_hoot):

Work = Force x displacement Force = mass x gravity Power = work / time

OpenStudy (irishboy123):

following @hoot_hoot 's lead, you might wish to start with the equations of motion gravity (g) exerts a constant force of 7g on the 7kg "something" over the displacement, and you need to know the displacement there are so many different formulations of the same thing, but you might consider \( x = ut + \frac{1}{2} g t^2 = \frac{1}{2} g t^2\) as it's *dropped* and not thrown..... so its initial velocity can be assumed to be zero here, \(x\) is the downward displacement power is more problematic. it is instantaneous, in terms of time. Power = work / time = Force x velocity